Calculate the mean for the following data set: 

Data Set = 2, 9, 10, 4, 8, 4, 12




1
Expert's answer
2021-03-19T09:49:06-0400
Add the numbers: sum = 2 + 9 + 10 + 4 + 8 + 4 + 12 = 49
Mean = sum / (total numbers) (there are 7 numbers): 49 / 7 = 7